# Calibration and Uncertainty for multiRater Volume Assessment in multiorgan Segmentation (CURVAS) challenge results

### v1

@@ 1 Introduction
Addressing these obstacles requires rigorous quality control, multi-expert annotations, and robust uncertainty quantification methods. To this end, the Calibration and Uncertainty for Multi-Rater Volume Assessment Challenge (CURVAS) challenge benchmarks algorithm performance on segmentation tasks using abdominal computed tomography (CT) scans with multi-expert annotations. CURVAS, organized as part of the Medical Image Computing and Computer-Assisted Intervention (MICCAI) 2024 conference, provides a structured environment for evaluating segmentation algorithms. Participants used a publicly available dataset[16] specifically constructed for this competition, featuring annotations from three expert radiologists who segmented three key abdominal organs: the pancreas, kidneys, and liver. This paper aims to provide an objective comparison of algorithms through quantitative and qualitative metrics, emphasizing inter-rater variability, calibration, and uncertainty quantification, beyond accuracy. Furthermore, CURVAS also assessed the ability of segmentation models to produce meaningful organ volume estimates, a clinically relevant metric, further highlights model strengths and limitations. A comprehensive quantitative and qualitative analysis reveals that the most accurate in segmenting are also the best calibrated models. However, models trained with pretrained networks or public datasets demonstrated greater robustness to distribution shifts, effectively detecting out-of-distribution regions. These insights contribute to the development of trustworthy AI-assisted medical imaging tools, promoting safer and more effective clinical adoption.
@@ 3.7 PedictED
BreizhSeg TS fine-tuned 20 times using four expert-specific training sets and one combined set3 Adaptable Bayesian NN & TS $128\!\times\!128\!\times\!$ 128 × 128
@@
DLAI All annotations used separately xLSTM-UNet adding a CAT* block within the encoder $128\!\times\!128\!\times\!$ 128 × 128
@@
BCNAIM Model per annotator UNet $96\!\times\!96\!\times\!$ 96 × 96
@@
CAI4CAI All annotations used separately MONAI: SegResNet $144\!\times\!144\!\times\!$ 144 × 144
@@
* The Spacing is specified below the Patch Size ** The Learning Rate is specified below the Optimizer *** GPU used for training, for inferences all teams had the same environments (1) with momentum of 0.99 & weight decay of 0.001 (2) RandCropByPosNegLabel crops fixed-size regions centered on a foreground or background voxel (pos=2, neg=1); RandRotate90 applies random 90° rotations (max k=3), and RandShiftIntensity shifts intensity within a 0.10 range. (3) each set followed a five-fold cross-validation strategy (4 sets × 5 folds = 20 runs) (4) with the weight decay $1\times 10^{-5}$ Linear warmup from $lr=0.0$ to 0.0001 for 100 epochs, then cosine annealing decay for the rest of the training
@@ 3.8 Teams’ summary
Out of the seven teams, three (MedIG, BreizhSeg, PredictED) submitted two algorithms, meaning the total number of submissions was ten. The different teams’ algorithms are summarized in Table 1. All of the submissions were based on a UNet architecture, each with its own variations. Two of the teams, MedIG and PredictED, incorporated merging techniques for the manual annotations provided, while the others opted not to use any merging and instead trained on all labels. All participants trained with small 3D patches, and during inference, they employed the sliding window technique with specific adaptations for each model, except for one, which trained using 2D images from entire CT slices. Most of the algorithms utilized data augmentation, and the majority combined Dice Loss and Cross Entropy for training. However, PredictED used Focal Loss to address class imbalance and CAI4CAI used a combination of Dice Loss and Cross Entropy, as other teams, but added a calibration loss based on Average Calibration Error. Lastly, four teams (PrAEcision, BreizhSeg, BCNAIM, and PredictED) incorporated external knowledge, either through pretrained models or public external databases.
@@ 4.1.2 Overall Algorithm Performance
Algorithm DSC (%) Confidence (%) ECE (×$\times 10^{-3}$) CRPS ($cm^{3}$)
@@ 4.1.3 Analysis Per Clinical Group
The focus of group-wise performance analysis, however, is not just the absolute performance of each algorithm but the trends observed across groups and whether these trends remain consistent for all methods. A clear pattern emerges: segmentation quality (measured by DSC) decreases progressively from Group B to Group A, reaching its lowest values in Group C. This aligns with previous observations that Group C presents the greatest challenges, containing more complex anatomical structures and ambiguous cases. Similarly, volume estimation errors (CRPS) increase in the same order (Group B $\xrightarrow{}$ Group A $\xrightarrow{}$ Group C) mirroring the trend in DSC values.
@@ A B C A B C
Algorithm ECE (×$\times 10^{-3}$) CRPS ($cm^{3}$)
@@ 4.3 Statistical Analysis
To analyze the statistical dispersion and robustness of algorithm performance, we employed a bootstrapping procedure to estimate the confidence in metric means and assess the stability of algorithm rankings. Specifically, we performed 500 bootstrap iterations by resampling the test cases with replacement. For each iteration, the mean metric score per algorithm was computed, followed by ranking all algorithms accordingly. From these bootstrap-derived rankings, we calculated the mean rank, standard deviation (as a measure of rank stability), and median rank (to provide a robust central tendency measure) for each algorithm. Additionally, we computed 95% confidence intervals for the mean ranks using ±$\pm 1.96$ times the standard deviation, offering a statistical indication of ranking reliability. Algorithms were then sorted by their mean rank, with lower values indicating better performance.
@@ 5.2 Clinical applications
Understanding and analyzing variability is essential in medical imaging, where discrepancies in expert annotations or algorithmic outputs can significantly impact diagnosis and patient care. Acknowledging this uncertainty and incorporating multiple perspectives enhances both the accuracy and reliability of clinical decisions. Several clinical applications could benefit from the integration of variability studies, particularly in areas such as radiology, oncology, and pathology. For instance, in radiology, variability studies could help refine tumor segmentation techniques, allowing for more accurate measurements of tumor size and location, which are critical for treatment planning and monitoring. In oncology, understanding variability in imaging data could improve the assessment of treatment response, as different radiologists or automated systems may interpret changes in tumor morphology differently. By quantifying these variations, clinicians could better account for uncertainties when making decisions about patient management. In pathology, variability studies could be applied to the classification of tissue samples, ensuring that multiple experts or algorithms arrive at consistent results, reducing the risk of misclassification in diagnosing diseases such as cancer. Furthermore, integrating variability analysis into clinical decision support systems could enhance personalized medicine, helping tailor treatment strategies based on the degree of uncertainty in individual cases. Ultimately, applying these insights to clinical practice would lead to more robust, transparent, and reliable diagnostic processes, providing clinicians with the information needed to make better-informed, patient-centered decisions.
@@ 6 Conclusion and future directions
A major challenge for future CURVAS editions will be not only identifying new clinical applications for concepts such as variability, uncertainty, and calibration, but also determining how to effectively communicate these aspects to clinicians. Conveying the importance of these elements is essential to ensure diagnostic decision-making in a more comprehensive understanding of both data and model limitations. This involves moving beyond presenting only point predictions to also providing insights into prediction confidence and expert disagreement. In parallel, the development of segmentation models should prioritize generalizability and the generation of well-calibrated, uncertainty-aware predictions. To achieve this, future efforts should emphasize the integration of diverse datasets, use of pre-trained models, and adoption of advanced calibration techniques—alongside the inclusion of expert discrepancies in both training and evaluation pipelines—to build more trustworthy and clinically relevant AI systems.

### v2

@@ 1 Introduction
Unlike previous segmentation challenges that primarily focused on label fusion or ignored annotator disagreement during evaluation, CURVAS is the first to systematically benchmark both segmentation accuracy and calibration under multi-rater uncertainty, with a focus on preserving and leveraging inter-rater variability. Addressing these obstacles requires rigorous quality control, multi-expert annotations, and robust uncertainty quantification methods. To this end, the Calibration and Uncertainty for multiRater Volume assessment in multiorgan Segmentation challenge (CURVAS) benchmarks algorithm performance on segmentation tasks using abdominal computed tomography (CT) scans with multi-expert annotations. CURVAS, organized as part of the Medical Image Computing and Computer-Assisted Intervention (MICCAI) 2024 conference, provides a structured environment for evaluating segmentation algorithms. Participants used a publicly available dataset[16] specifically constructed for this competition, featuring annotations from three expert radiologists who segmented three key abdominal organs: the pancreas, kidneys, and liver.
CURVAS is distinguished not only by its evaluation of segmentation accuracy but also by its emphasis on calibration and uncertainty under multi-rater settings. This paper aims to provide an objective comparison of algorithms through quantitative and qualitative metrics, emphasizing inter-rater variability, calibration, and uncertainty quantification, beyond accuracy. Furthermore, the challenge also assessed the ability of segmentation models to produce meaningful organ volume estimates, a clinically relevant metric, further highlighting each model’s strengths and limitations. A comprehensive analysis reveals that the most accurate models in terms of segmentation are also those best calibrated. Additionally, models trained with pretrained networks or public datasets demonstrated greater robustness to distribution shifts, effectively identifying out-of-distribution regions. These insights contribute to the development of trustworthy AI-assisted medical imaging tools, supporting safer and more effective clinical adoption.
@@ 3.7 PedictED
BreizhSeg TS fine-tuned 20 times using four expert-specific training sets and one combined set3 Adaptable Bayesian NN & TS $128\!\times\!128\!\times\!128$
@@
DLAI All annotations used separately xLSTM-UNet adding a CAT* block within the encoder $128\!\times\!128\!\times\!128$
@@
BCNAIM Model per annotator UNet $96\!\times\!96\!\times\!96$
@@
CAI4CAI All annotations used separately MONAI: SegResNet $144\!\times\!144\!\times\!144$
@@
* The Spacing is specified below the Patch Size ** The Learning Rate is specified below the Optimizer *** GPU used for training, for inferences all teams had the same environments (1) with momentum of 0.99 & weight decay of 0.001 (2) RandCropByPosNegLabel crops fixed-size regions centered on a foreground or background voxel (pos=2, neg=1); RandRotate90 applies random 90° rotations (max k=3), and RandShiftIntensity shifts intensity within a 0.10 range. (3) each set followed a five-fold cross-validation strategy (4 sets × 5 folds = 20 runs) (4) with the weight decay $1\times 10^{-5}$ & Linear warmup from $lr=0.0$ to 0.0001 for 100 epochs, then cosine annealing decay for the rest of the training
@@ 3.8 Teams’ summary
Out of the seven teams, three (MedIG, BreizhSeg, PredictED) submitted two algorithms, meaning the total number of submissions was ten. The different teams’ algorithms are summarized in Table 1. All of the submissions were based on a UNet architecture, each with its own variations. Two of the teams, MedIG and PredictED, incorporated merging techniques for the manual annotations provided (using STAPLE and Majority Voting, respectively), while the others opted not to use any merging and instead trained on all labels. All participants trained with small 3D patches, and during inference, they employed the sliding window technique with specific adaptations for each model, except for one, which trained using 2D images from entire CT slices. Preprocessing strategies varied: MedIG, PrAEcision, and BCNAIM standardized voxel spacing and used intensity normalization; BCNAIM also performed cropping centered on organ regions, while CAI4CAI scaled intensities and applied affine transformations; PredictED worked on native 2D CT slices without resampling. Augmentation usage ranged from none (BreizhSeg) to highly diverse pipelines: DLAI combined spatial (rotations, scaling, mirroring) and intensity (noise, brightness, contrast, gamma) transformations with elastic deformations; MedIG used nnU-Net’s default 3D augmentations with mirroring disabled; PrAEcision applied spatial transformations; BCNAIM used MONAI transforms targeting both spatial and intensity variability; PredictED employed 2D geometric (rotations, flips, shifts, scales) and photometric (brightness, contrast) augmentations. Most algorithms combined Dice Loss and Cross Entropy for training. However, PredictED used Focal Loss to address class imbalance and CAI4CAI used a combination of Dice Loss and Cross Entropy, as other teams, but added a calibration loss based on Average Calibration Error. Lastly, four teams (PrAEcision, BreizhSeg, BCNAIM, and PredictED) incorporated external knowledge, either through pretrained models or public external databases.
@@ 4.1.2 Overall Algorithm Performance
Algorithm DSC (%) Confidence (%) ECE ($\times 10^{-3}$) CRPS ($cm^{3}$)
@@ 4.1.3 Analysis Per Clinical Group
[…]
@@ A B C A B C
Algorithm ECE ($\times 10^{-3}$) CRPS ($cm^{3}$)
@@ 4.3 Statistical Analysis
The statistical analysis of the results comprises two complementary components: first, an assessment of the statistical significance of performance differences between algorithms; and second, an evaluation of the stability and robustness of the resulting rankings.
Figure 6 presents the pairwise statistical significance results based on two-sided Wilcoxon signed-rank tests conducted across all participating methods. The comparisons were performed separately for each evaluation metric, Dice Score (DSC), Expected Calibration Error (ECE), Continuous Ranked Probability Score (CRPS), and confidence reliability, with p-values adjusted using Bonferroni correction to account for multiple testing. In the heatmaps, darker shades correspond to smaller adjusted p-values, indicating statistically significant differences between method pairs.
This analysis serves to reinforce and contextualize performance-based rankings by identifying which observed differences are statistically robust. Notably, the top-performing methods, MedIG, PrAEcision, and BreizhSeg, consistently exhibit statistically significant improvements in calibration and uncertainty-related metrics (ECE, CRPS, and confidence reliability) when compared to lower-ranking teams such as PredictED, CAI4CAI, and BCNAIM. In contrast, the Dice Score matrix reveals fewer statistically significant differences, underscoring that segmentation accuracy alone does not fully capture model quality, highlighting the importance of comprehensive evaluation in multi-rater, uncertainty-aware settings.
Figure 6: Pairwise statistical comparisons (corrected p-values) across segmentation algorithms for four performance metrics: Dice Score (DSC), Confidence, Expected Calibration Error (ECE), and Continuous Ranked Probability Score (CRPS). Each heatmap shows the results of pairwise Wilcoxon signed-rank tests between methods, with p-values corrected for multiple comparisons using the Holm–Bonferroni method. Darker cells (blue) indicate statistically significant differences ($p<0.05$), while lighter cells indicate non-significant results after correction. Diagonal values are 1.0 by definition (self-comparisons).
To analyze the statistical dispersion and robustness of algorithm performance, we employed a bootstrapping procedure to estimate the confidence in metric means and assess the stability of algorithm rankings. Specifically, we performed 500 bootstrap iterations by resampling the test cases with replacement. For each iteration, the mean metric score per algorithm was computed, followed by ranking all algorithms accordingly. From these bootstrap-derived rankings, we calculated the mean rank, standard deviation (as a measure of rank stability), and median rank (to provide a robust central tendency measure) for each algorithm. Additionally, we computed 95% confidence intervals for the mean ranks using $\pm 1.96$ times the standard deviation, offering a statistical indication of ranking reliability. Algorithms were then sorted by their mean rank, with lower values indicating better performance.
@@
It is worth noting that the plotted bubbles represent the mean aggregated rank for each method, whereas the 95% confidence intervals are obtained by bootstrapping per-case scores before rank aggregation. Because the rank transformation is nonlinear, the mean of the bootstrapped ranks does not necessarily match the mean aggregated rank. This can lead to apparent misalignments between the bubble centers and their corresponding confidence intervals, which reflect uncertainty in the underlying per-case scores rather than in the aggregated ranks.
@@ 5.1 Best Performing methods
We have used pairwise Wilcoxon signed-rank tests across metrics to study whether there are statistically significant differences between the performances of the methods. We found that, based on the corrected p-value matrices (see Figure 6), there is no statistically significant difference between the top-performing methods, MedIG, PrAEcision, and BreizhSeg, across any of the four evaluation metrics, suggesting their performance is comparable and robust with respect to both segmentation accuracy and uncertainty calibration. However, performance differences between teams translated into statistically significant improvements, particularly in terms of uncertainty calibration. We observed that top-performing methods consistently outperformed lower-ranking teams such as PredictED, BCNAIM, DLAI, and CAI4CAI ($p<0.05$ after Bonferroni correction).
Notably, these significance tests revealed that differences in Dice score were less pronounced, suggesting that segmentation accuracy alone does not capture the full spectrum of model quality. Instead, calibration metrics better reflected performance differences among the top contenders, reinforcing the importance of evaluating not just what the model predicts, but how confident it is in those predictions. This confirms that uncertainty-aware models can meaningfully outperform others even when segmentation accuracy appears comparable which shows the necessity of a holistic evaluation strategy that goes beyond pixel-wise overlap to assess trustworthiness and calibration in clinical AI systems. This statistical analysis complements the performance-based rankings by identifying where observed differences are robust and reproducible.
@@ 5.2 Clinical applications
Understanding and analyzing variability is essential in medical imaging, where discrepancies in expert annotations or algorithmic outputs can significantly impact diagnosis and patient care. Acknowledging this uncertainty and incorporating multiple perspectives enhances both the accuracy and reliability of clinical decisions. Several clinical applications could benefit from the integration of variability studies, particularly in areas such as radiology, oncology, and pathology. For instance, in oncology, understanding variability in imaging data could improve the assessment of treatment response, as different radiologists or automated systems may interpret changes in tumor morphology differently. In radiology, variability studies could help refine tumor segmentation techniques, allowing for more accurate measurements of tumor size and location, which are critical for treatment planning and monitoring. For example, when radiologists assess vascular invasion in the pancreas—a critical factor for determining tumor resectability—the evaluation heavily relies on clearly defining the tumor’s boundaries. However, these boundaries often vary between experts due to image ambiguity and low contrast. A calibrated model could flag regions with high uncertainty, such as the tumor margins near vessels, and explicitly alert the clinician. This would allow radiologists to consider not only the model’s prediction but also the reliability of that prediction, enabling more nuanced interpretations. Ultimately, integrating such uncertainty-aware visual cues into diagnostic workflows could improve both trust in AI systems and decision quality in borderline or ambiguous cases.By quantifying these variations, clinicians could better account for uncertainties when making decisions about patient management. In pathology, variability studies could be applied to the classification of tissue samples, ensuring that multiple experts or algorithms arrive at consistent results, reducing the risk of misclassification in diagnosing diseases such as cancer. Furthermore, integrating variability analysis into clinical decision support systems could enhance personalized medicine, helping tailor treatment strategies based on the degree of uncertainty in individual cases. In the end, applying these insights to clinical practice would lead to more robust, transparent, and reliable diagnostic processes, providing clinicians with the information needed to make better-informed, patient-centered decisions.
@@ 6 Conclusion and future directions
A major challenge for future CURVAS editions will be not only identifying new clinical applications for concepts such as variability, uncertainty, and calibration, but also determining how to effectively communicate these aspects to clinicians. Conveying the importance of these elements is essential to ensure diagnostic decision-making in a more comprehensive understanding of both data and model limitations. This involves moving beyond presenting only point predictions to also providing insights into prediction confidence and expert disagreement. In parallel, the development of segmentation models should prioritize generalizability and the generation of well-calibrated, uncertainty-aware predictions. While our dataset provides a valuable benchmark, its size and source may limit generalizability across institutions or scanner types. Future work should explore strategies to mitigate domain shift, such as incorporating domain adaptation techniques or fine-tuning models on site-specific data. To achieve this, future efforts should emphasize the integration of diverse datasets as well, use of pre-trained models, and adoption of advanced calibration techniques—alongside the inclusion of expert discrepancies in both training and evaluation pipelines—to build more trustworthy and clinically relevant AI systems.